OpenStudy (anonymous):

solve using square root method 9x^2-25=0

OpenStudy (amistre64):

9x^2 = 25 ...add 25 to each side sqrt(9x^2) = sqrt(25) ... square root each side 3x = 5 x= 5/3 or x=(-5/3)